# Perpetual Future Contracts in Centralized and Decentralized Exchanges: Mechanism and Traders’ Behavior

### v2

@@ Abstract
This study presents a groundbreaking Systematization of Knowledge (SoK) initiative, focusing on an in-depth exploration of the dynamics and behavior of traders on perpetual future contracts across both centralized exchanges (CEXs) and decentralized exchanges (DEXs). We summarize the features of CEXs and DEXs and propose 4 exchange models that govern the operations of exchanges in different type, allowing for the classification of any exchange into one of these predefined models Our research also includes an empirical analysis of traders behavior in exchanges of the 4 types of models. On DEX of Virtual Automated Market Making (VAMM) Model, open interest on short and long positions exert effect on price volatility in opposite direction, attributable to VAMM’s price formation mechanism. In the DEXs with Oracle Pricing Model, we observed a distinct asymmetry in trader behavior between buyers and sellers. Such asymmetry might stem from uninformed traders reacting more strongly to positive news than to negative, leading to a tendency to accumulate long positions. This study sheds light on the potential risks and advantages of using perpetual future contracts within the DeFi space while provides mathematical basis and empirical insights based on which future theoretical works can be configurated, offering crucial insights into the rapidly evolving world of blockchain-based financial instruments.
Keywords Blockchain  $\cdot$ Perpetual Futures  $\cdot$ Volatility  $\cdot$ DeFi.
@@ 1 Introduction
Over the past decade, blockchain technology has undergone swift and substantial development, garnering an extensive user base exceeding 200 million individuals on a global scale. Among those applications, cryptocurrency derivative tradings are implemented among exchanges organized in both centralized and decentralized manner, within which perpetual futures came to the most prominence with daily trading volume of more than 100 billion  [^1] At the time of writing, 22 November 2023, the 24H trading volume of perpetual future contracts among all the exchanges is \$109.9 billion (Source: https://coinalyze.net).. Investors prefer perpetual future contracts to future contracts, reflected by the far less 24H volume of future contracts (\$1.1 billion), mainly because of the efficiency of perpetuals in hedging and speculation with high leverage while without the need of delivery and rollover. Before 2022, investors trade perpetual futures on centralized exchanges (CEXs), but following the collapse of FTX there emerged many decentralized exchanges (DEXs) carrying perpetual futures, which are operated with more elements on-chain. Since the DEXs are constructed with smart contracts involved, more transparency can be ensured, helping guard against the fraudulence incurred in CEXs.
Although both running perpetual futures, CEXs and DEXs are considerably different in terms of the order matching, price discovering, and custody of investors’ fund, leading to different traders’ behavior and implications on the market micro-structure. Unlike the case of CEXs, which adopt the limit-order book to form prices, orders submitted by traders in DEXs are matched with liquidity providers and the prices are formed based on the relative abundance of assets in the liquidity pool. The great difference in the design of market making and price discovering discourages us from using traditional micro-structure model, e.g., Kyle’s model (1985) [1], to understand DEXs. Therefore, in this paper we ask, what are the fundamental differences of the design among DEXs and CEXs carrying perpetual futures, how those differences affect price formation and investors’ behavior, and what kinds of empirical implications can be provided to the potential theoretical models for the micro-structure of DEXs? To this end, in the first part of this study we delineate a detailed Systematization of Knowledge (SoK) contrasting the operational paradigms of Centralized Exchanges (CEXs) and Decentralized Exchanges (DEXs) in the context of perpetual futures trading. We summarize the features of CEXs and DEXs and propose 4 exchange models that govern the operations of exchanges in different type, allowing for the classification of any exchange into one of these predefined models. In the second part, we examine traders’ behavior in exchanges under different exchange models with data from August 2020 to September 2023, revealing distinctive relationship between price formation and trader behavior derived from the fundamental design differences among different exchange models. This study contributes to the literature by providing a classification standard for the design of DEXs and CEXs on which perpetual future contracts are traded, while elaborating the distinctive features of each exchange model. The trading process and participants’ payoffs are formalized separately for each exchange model, offering basis on which future works can be constructed. The pattern revealed by our empirical study corroborates the differences implied by different exchange model, constituting the first sort of empirical research on how traders behave on DEXs and shading lights on potential theoretical frameworks for DEXs.
We categorize CEXs and DEXs based on their Custody Module, with CEXs (CEX Model) managing traders’ funds off-chain and DEXs (DEX models) recording and managing funds on-chain via smart contracts. DEXs are further classified into three exchange models: Hybrid, Oracle Pricing, and Virtual Market Making (VAMM). The Hybrid Model, straddling CEX Model and the other DEX models, still employs limit-order books for trade matching, while the Oracle Pricing Model and VAMM Models utilize smart contracts. In the latter two, participants are categorized into active traders (Traders) and passive liquidity providers (Liquidity Providers), the latter accepting orders and earning transaction fees for their liquidity provision. The decentralized, transparent nature of DEXs, underpinned by smart contracts, mitigates risks of misconduct and fraud, highlighting their advantages.
@@
We investigate the distinct behaviors of traders on exchanges within different models by examining the relationship between the price volatility of the underlying asset, i.e., Bitcoin, and various trading activities, i.e., trading volume, open interest, liquidation, and leverage. Our study encompasses prominent CEXs such as Binance and Bybit, alongside DEXs like GMX, GNS, and Perpetual Protocol V2, with the latter two operating under the Oracle Pricing Model and Perpetual Protocol V2 employing the Virtual Automated Market Making (VAMM) Model. The empirical evidence shows that, for CEXs, price volatility is positively related with trading volumes but negatively related with open interests, consistent with findings in the traditional future markets. While the effects of trading activity on price volatility in CEXs reflect the market depth and can be explained by Kyle’s model [1], the VAMM Model introduces a nuanced dynamic in the impact of open interest, varying between long and short positions. This asymmetry is attributable to the VAMM’s price formation mechanism, where the rate of change in an asset’s relative price inversely correlates with its abundance in the liquidity pool. Consequently, market depth increases with rising open interests in short positions, as the underlying asset accumulates in the liquidity pool, and the reverse holds true for long positions. While in CEXs and exchange of VAMM Model traders’ trading activity help form the price, traders in exchanges of Oracle Pricing Model (GMX and GNS) accept the price offered by the Oracle and thus act as pure price takers. Therefore, trading activity should be interpreted as purely traders’ reaction to the price change of the underlying asset. When the price become more volatile, trading volumes increase while open interests decrease, with change in long and short positions different. These empirical estimations can be explained by the predictions based on Shalen’s dispersion of beliefs model (1993) [2], which address the asymmetry of information traders can access. We also find that uninformed traders tend to overreact more to positive news than negative, evidenced by more of the long positions accumulated.
This comprehensive examination not only delineates the distinct behavioral patterns of traders across exchanges of different design (CEXs and DEXs) but also provides a deeper understanding of the underlying mechanisms driving these behaviors, offering valuable insights for future research and practical applications in the evolving landscape of digital asset trading. While the primary focus of our research is on Decentralized Exchanges (DEXs) operating with perpetual contracts, the findings and theoretical frameworks developed herein have broader applicability, extending to other trading paradigms within DEXs that exhibit the fundamental bifurcation between Traders and Liquidity Providers. This is exemplified in platforms such as PancakeSwap (https://pancakeswap.finance) and Uniswap (https://uniswap.org). Furthermore, as the Decentralized Finance (DeFi) ecosystem evolves, the proliferation of synthetic assets is anticipated to enhance the diversification of portfolios managed on DEXs. The insights gleaned from this study are thus of paramount importance to both academic researchers and industry practitioners.
This paper is organized into distinct sections for clarity and depth of exploration. Section 2 provides a discussion of literature related to our work. Section 3 delineates the foundational principles underpinning perpetual trading within both the realms of Decentralized Finance (DeFi) and Centralized Finance (CeFi). In Section 4, we present a systematic and exhaustive survey of the various models associated with perpetual futures, framed within a Systematization of Knowledge (SoK) approach. Section 5 embarks on an empirical investigation into the trading behaviors exhibited on five prominent perpetual futures platforms: GMX, GNS, Perpetual Protocol V2, Binance, and Bybit. The paper culminates with Section 6, where we offer concluding observations and reflections on the implications of our findings for the broader field of blockchain-based financial instruments and their role in the evolving landscape of finance.
2 Related Literature
In board terms, our study relates to three sorts of literature: those examining Defi, cryptocurrency future contracts, and Automated Market Making (AMM). The literature on Defi includes works discussing the fragility and inefficiency of the Defi lending platforms (e.g., [3, 4, 5]), those examining the liquidation and leverage risks (e.g.,[6, 7, 8, 9, 10, 11, 12, 13]), and the ones comparing Defi and Cefi platforms (e.g., [14, 15]). Qin et al. [10] investigated four major Defi protocols involving lending (MakerDAO [^2] https://makerdao.com/., dYdX [^3] https://dydx.exchange/., Aave [^4] https://aave.com/., and Compound [^5] https://compound.finance/.), revealing that liquidators can optimize strategies to maximize profits, potentially increasing borrowers’ losses during liquidations. It underscores the inherent risks and instabilities in DeFi markets, particularly through the lens of liquidation mechanics and their impact on market dynamics. Later, Wang et al. [16] analyzed the risks associated with on-chain leverage in the DeFi sector by formalising a model for under-collateralized DeFi lending platforms. They provide evidence that high levels of leverage can lead to significant risks and instabilities. Since liquidation risk is also considered as one of the major risks of perpetual futures, this study evaluates differences in traders’ behavior in CEXs and DEXs revealed by liquidation.
@@ 2 Related Literature
Existing literature evaluates cryptocurrency future contracts mainly in terms of their relationship with the spot market. Akyildirim et al. [17] studied the impact of Bitcoin futures on the cryptocurrency market, particularly the introduction of CME and CBOE futures contracts in December 2017. Alexander et al. [18] found that BitMEX derivatives lead the price discovery process over major Bitcoin spot exchanges. Hung et al. [19] identifies substantial pricing effects and breakpoints in market efficiency, indicating the dominant role of Bitcoin futures in price discovery compared to spot markets. As a special case of future contracts, perpetual future contracts are scarcely investigated, although with much higher trading volume. Besides the theoretical discussion on the arbitrage between perpetual future markets and spot markets [20], there lacks enough empirical works examining perpetual future traders’ behavior, especially in DEXs. After the pioneering work by Soska et al. [21], which conducted the first analysis on the investor profile for perpetual future contracts in BitMEX, i.e., a CEX, Alexander et al. [22] constructed the optimal hedging strategy with empirical corroboration.

### v4

# Exploring the Impact: How Decentralized Exchange Designs Shape Traders’ Behavior on Perpetual Future Contracts

@@ Abstract
In this paper, we analyze traders’ behavior within both centralized exchanges (CEXs) and decentralized exchanges (DEXs), focusing on the volatility of Bitcoin prices and the trading activity of investors engaged in perpetual future contracts. We categorize the architecture of perpetual future exchanges into three distinct models, each exhibiting unique patterns of trader behavior in relation to trading volume, open interest, liquidation, and leverage. Our detailed examination of DEXs, especially those utilizing the Virtual Automated Market Making (VAMM) Model, uncovers a differential impact of open interest on long versus short positions. In exchanges which operate under the Oracle Pricing Model, we find that traders primarily act as price takers, with their trading actions reflecting direct responses to price movements of the underlying assets. Furthermore, our research highlights a significant propensity among less informed traders to overreact to positive news, as demonstrated by an increase in long positions. This study contributes to the understanding of market dynamics in digital asset exchanges, offering insights into the behavioral finance for future innovation of decentralized finance.
keywords:
Blockchain. Perpetual Futures. Volatility. DeFi.
@@ 1 Introduction
Over the past decade, blockchain technology has seen rapid development, attracting over 200 million users globally. Among its applications, cryptocurrency derivative trading, particularly perpetual futures, has become prominent, with daily trading volumes exceeding \$100 billion. Perpetual futures, introduced into the cryptocurrency market by BitMEX (https://bitmex.com/) in 2016 and conceptualized by Robert J. Shiller[19], represent a novel variant of traditional futures contracts. Unlike their traditional counterparts, perpetual futures do not have a fixed delivery date with their prices pegged to spot prices of the underlying assets, allowing for indefinite trading without the need for contract rollovers.Investors show a preference for perpetual futures over traditional futures due to their efficiency in hedging and speculation, high leverage, and the absence of delivery and rollover requirements. Before 2022, investors trade perpetual futures on centralized exchanges (CEXs), but following the collapse of FTX there emerged many decentralized exchanges (DEXs) carrying perpetual futures, which are operated with more elements on-chain. Since the DEXs are constructed with smart contracts involved, more transparency can be ensured, helping guard against the fraudulence incurred in CEXs.
Although both running perpetual futures, CEXs and DEXs are considerably different in terms of their order matching and price discovering, leading to different traders’ behavior and implications on the market micro-structure. The design of the current CEXs and DEXs can be classified into three exchange models according to the way orders are matched: Limit-Order-Book (LOB) Model, Virtual Market Making (VAMM) Model, and Oracle Pricing Model. Unlike the case of LOB Model, which adopt the limit-order book to form prices, orders submitted by traders in exchanges of VAMM Model are matched with liquidity providers  [^1] Liquidity providers deposit assets in liquidity pools facilitated by smart contracts, where traders execute transactions. and the prices are formed based on the relative abundance of assets in the liquidity pool. While in exchanges of LOB and VAMM Model traders’ trading activity help form the price, traders in exchanges of Oracle Pricing Model accept the price offered by the Oracle and thus act as pure price takers. Although liquidity providers are still passively act as the trading counterparty, the pricing is governed by a Price Oracle, which assimilates prices from various spot exchanges, thereby rendering transactions in exchanges of Oracle Pricing Model incapable of impacting the price directly [^2] Details about the design of different exchange models can be found in Appendix A or the Systems of Knowledge paper [8].. Therefore, in this paper we ask how those differences in exchanges of difference exchange models affect price formation and investors’ behavior and what kinds of empirical implications can be provided to the potential theoretical models for those exchange models.
@@
We investigate the distinct behaviors of traders on exchanges within different models by examining the relationship between the price volatility of the underlying asset, i.e., Bitcoin, and various trading activities, i.e., trading volume, open interest, liquidation, and leverage. Our study encompasses prominent exchanges of LOB Model such as Binance (https://www.binance.com/en) and Bybit (https://www.bybit.com), alongside exchanges of Oracle Pricing Model like GMX (https://gmx.io) and GNS (https://gains.trade), and Perpetual Protocol V2 (https://perp.com) which employs VAMM Model. The empirical evidence shows that, for exchanges of LOB Model (Binance and Bybit), price volatility is positively related with trading volumes but negatively related with open interests, consistent with findings in the traditional future markets. While the effects of trading activity on price volatility in exchanges of LOB Model reflect the market depth and can be explained by Kyle’s model [13], the VAMM Model (Perpetual Protocol V2) introduces a nuanced dynamic in the impact of open interest, varying between long and short positions. This asymmetry is attributable to the VAMM’s price formation mechanism, where the rate of change in an asset’s relative price inversely correlates with its abundance in the liquidity pool. Consequently, market depth increases with rising open interests in short positions, as the underlying asset accumulates in the liquidity pool, and the reverse holds true for long positions. While in exchanges of LOB and VAMM Model traders’ trading activity help form the price, traders in exchanges of Oracle Pricing Model (GMX and GNS) accept the price offered by the Oracle and thus act as pure price takers. Therefore, trading activity should be interpreted as purely traders’ reaction to the price change of the underlying asset. When the price become more volatile, trading volumes increase while open interests decrease, with change in long and short positions different. These empirical estimations can be explained by the predictions based on Shalen’s dispersion of beliefs model (1993) [18], which address the asymmetry of information traders can access. We also find that uninformed traders tend to overreact more to positive news than negative, evidenced by more of the long positions accumulated.
@@
Existing literature evaluates cryptocurrency future contracts mainly in terms of their relationship with the spot market. Akyildirim et al. [1] studied the impact of Bitcoin futures on the cryptocurrency market, especially the introduction of CME and CBOE futures contracts in December 2017. Alexander et al. [2] found that BitMEX derivatives lead the price discovery process over Bitcoin spot markets. Hung et al. [12] identifies substantial pricing effects and breakpoints in market efficiency, indicating the dominant role of Bitcoin futures in price discovery compared to spot markets. As a special case of future contracts, perpetual future contracts are scarcely investigated, although with much higher trading volume. Besides the theoretical discussion on the arbitrage between perpetual future markets and spot markets [11], there lacks enough empirical works examining perpetual future traders’ behavior, especially in DEXs. After the pioneering work by Soska et al. [20], which conducted the first analysis on the trader profile for perpetual future contracts in BitMEX, i.e., a CEX, Alexander et al. [4] constructed the optimal hedging strategy with empirical corroboration.